OpenStudy (anonymous):

Write the equation of the line that is perpendicular to the line y = 2x + 2 and passes through the point (6, 3).

OpenStudy (skullpatrol):

Any ideas?

OpenStudy (mertsj):

What is the slope of the given line?

OpenStudy (anonymous):

perpendicular line would have the negative reciprocal slope i believe, that is -1/2

OpenStudy (anonymous):

y=ax+b where y=3, x=6, a=-1/2, so now solve for b and you're good

OpenStudy (anonymous):

can i just get the answer please i need to have this done

OpenStudy (anonymous):

no, try to work through it. this is not a homework answers site. we are here to help YOU solve it.

OpenStudy (anonymous):

pi*6

OpenStudy (anonymous):

i had B

OpenStudy (anonymous):

y=(1/2)x+(1/10)